Jefferson Lab HALL A Logbook for August 09 - All EXCEPT Automatic Entries

  • Main INDEX

    Monday

  • 287278 08/31/09 23:58 Aniol shift summary
  • 287275 08/31/09 23:41 M Friend Compton Photon Detector Pedestal Shift
  • 287274 08/31/09 23:40 dalton Bug in pan?
  • 287273 08/31/09 23:29 Aniol target screen
  • 287272 08/31/09 23:28 Aniol target graphs
  • 287271 08/31/09 23:15 dalton conf/control.conf_reg altered
  • 287270 08/31/09 23:02 O.Glamazdin The Moller spin dance results
  • 287263 08/31/09 22:28 aniol raster pattern
  • 287261 08/31/09 22:25 M Friend FADC Compton Spectrum
  • 287258 08/31/09 22:04 Aniol finished Moller at 21:45
  • 287257 08/31/09 22:04 Aniol target: loop 3 in beam position
  • 287256 08/31/09 21:52 mercado Moller measurements done
  • 287255 08/31/09 21:32 mercado Updated PAN in ~/happex3/pan
  • 287254 08/31/09 19:13 mcnulty Q2 from EKRQ2 looks good
  • 287253 08/31/09 19:09 R. Michaels R-HRS Qsq
  • 287252 08/31/09 18:53 adaq We are currently doing moller measurements.
  • 287251 08/31/09 18:17 mercado ATTENTION Shift Workers! Online Monitoring using PANGUIN
  • 287249 08/31/09 16:58 R. Michaels ADC18 DAC noise glitch, etc.
  • 287247 08/31/09 16:42 R. Michaels re: ADCX6 (ROC26, 4th slot)
  • 287243 08/31/09 16:04 jpchen Day shift summary
  • 287241 08/31/09 15:37 mercado Changed datamap for HAPPEX detectors
  • 287240 08/31/09 15:20 mercado Another HAPPEX DAQ Change
  • 287238 08/31/09 15:17 W. Deconinck Target screen capture
  • 287237 08/31/09 15:17 jpchen tools screen
  • 287234 08/31/09 14:46 A. Camsonne GSO HV lowered to 1650 V
  • 287231 08/31/09 14:35 R. Michaels tried, then backed out, new CRL
  • 287228 08/31/09 14:31 A. Camsonne Compton electron detector in beam position
  • 287227 08/31/09 14:25 silwal happex I (H-C) base swapped with happex II (H-3)
  • 287226 08/31/09 14:17 silwal pedestal rms width fixed
  • 287223 08/31/09 13:49 jpchen In Beam Permit now
  • 287222 08/31/09 13:30 M Friend Compton FADC DAQ Change
  • 287220 08/31/09 13:22 A. Camsonne Compton electron noise analog board
  • 287208 08/31/09 12:42 A. Camsonne Compton electron detector in garage position
  • 287207 08/31/09 12:22 A. Camsonne ETROC inputs
  • 287204 08/31/09 11:33 R. Michaels re: ADC18 in R-HRS
  • 287203 08/31/09 11:32 A. Camsonne Sending 40 MHz clock to ETROC
  • 287202 08/31/09 10:59 R. Michaels runbird -> runbird.luis2
  • 287199 08/31/09 10:49 R. Michaels DAQ software changes
  • 287195 08/31/09 08:44 A. Camsonne Fixed left Compton electron detector
  • 287193 08/31/09 08:11 R. Subedi access activity (host expt may ignore it)
  • 287192 08/31/09 07:16 dalton Shift summary
  • 287191 08/31/09 06:29 dalton CHL is up, 1 more hour for recovery
  • 287190 08/31/09 05:04 paschke charge feedback working
  • 287189 08/31/09 04:55 paschke dithering control system working
  • 287188 08/31/09 04:50 paschke pedestal problems when using the 16-bit boards
  • 287187 08/31/09 04:23 mercado HAPPEX DAQ changes
  • 287174 08/31/09 02:00 dalton Cryo says 4 hours down
  • 287173 08/31/09 01:24 dalton Beam down for 2 K cold box problem
  • 287172 08/31/09 01:22 dwang Target moved to empty
  • 287164 08/31/09 00:20 silwal R/L det pedstal changed
  • 287161 08/31/09 00:01 rsholmes Owl shift summary
  • Sunday

  • 287155 08/30/09 23:31 silwal 60 uA current
  • 287147 08/30/09 22:55 A. Camsonne Compton laser power reduced
  • 287146 08/30/09 22:32 Kalyan Run 11802 is junk
  • 287143 08/30/09 22:23 A. Camsonne Electron detector Compton rates after moving detector back in position
  • 287142 08/30/09 22:18 Abdurahim Target main window snapshot
  • 287141 08/30/09 22:17 Abdurahim target heater strip chart
  • 287140 08/30/09 22:14 Abdurahim target strip charts
  • 287135 08/30/09 21:47 A. Camsonne Compton electron rate gone
  • 287134 08/30/09 21:16 A. Camsonne Standalone run EDet 18534
  • 287131 08/30/09 21:10 M Friend Compton Photon Detector Change
  • 287130 08/30/09 20:54 Abdurahim Heater alarm due to a beam trip
  • 287129 08/30/09 20:25 Abdurahim Target Main Screen snap shot
  • 287128 08/30/09 20:24 Abdurahim Target Heater Strip Chart
  • 287127 08/30/09 20:22 Abdurahim Target Strip Charts
  • 287126 08/30/09 20:22 Kalyan prompt analysis: Run 11795, beam modulation slopes are empty
  • 287123 08/30/09 20:11 rsholmes My bad
  • 287120 08/30/09 19:16 A. Camsonne Last version of ETROC firmware loaded
  • 287115 08/30/09 17:59 Kalyan prompt analysis not working
  • 287114 08/30/09 17:46 Abdurahim Target Screen
  • 287113 08/30/09 17:44 Abdurahim Target heater L3 strip chart
  • 287112 08/30/09 17:43 Abdurahim Target Strip Charts
  • 287111 08/30/09 17:40 Abdurahim Target
  • 287108 08/30/09 17:28 M Friend Compton Photon Detector HV Trip
  • 287107 08/30/09 16:44 Abdurahim Target
  • 287104 08/30/09 16:08 target opera day shift
  • 287103 08/30/09 16:04 silwal day shift summary
  • 287097 08/30/09 14:57 mcnulty Preliminary look at Rarm Q2 from run 4382
  • 287091 08/30/09 13:45 target opera beam trip L3 heater
  • 287088 08/30/09 13:27 dalton Prompt analysis
  • 287087 08/30/09 12:58 Kai Pan onlana.odef modified for getting the "Unser" in rootfile
  • 287086 08/30/09 12:51 silwal adc15 (bcm1/2) high gain, adc saturated
  • 287079 08/30/09 12:13 target opera L3 heater beam trip
  • 287077 08/30/09 11:58 mcnulty Larm alignment
  • 287076 08/30/09 11:54 target opera target empty L3 H
  • 287075 08/30/09 11:35 target opera target L3 H empty
  • 287069 08/30/09 10:13 silwal Det HV values
  • 287068 08/30/09 10:11 silwal Right arm HV turned on
  • 287065 08/30/09 09:55 Jack (Techs) Trip point for cryogen levels on LQ3
  • 287064 08/30/09 09:54 silwal magnetistrips after LQ3 reset
  • 287063 08/30/09 09:52 silwal magnet status
  • 287062 08/30/09 09:50 silwal L Q3 magnet up
  • 287060 08/30/09 09:33 silwal beam back up at 9:32 am, left arm, Q3 still down
  • 287056 08/30/09 08:50 target opera temperature graph
  • 287055 08/30/09 08:49 target opera power pressure graphs
  • 287054 08/30/09 08:48 mcnulty Larm HapDet Alignment
  • 287053 08/30/09 08:41 buddhini cryogenics level, spectrometer magnets.
  • 287052 08/30/09 08:37 buddhini cryogenics levels, magnet
  • 287049 08/30/09 08:21 mercado Fan Speed and Raster Size Change Summary
  • 287048 08/30/09 08:12 Target Opera loop pressure
  • 287047 08/30/09 08:02 mercado Boiling Studies - Ramping Current
  • 287046 08/30/09 07:47 ZAFAR AHMED Owl Shift Summary
  • 287029 08/30/09 06:18 A. Camsonne Compton electron rates with threshold
  • 287028 08/30/09 06:12 A. Camsonne Run single Compton Electron detector
  • 287027 08/30/09 06:08 mercado Switched HAPPEX DAQ Configuration
  • 287008 08/30/09 05:13 pbaturin correction
  • 287006 08/30/09 05:10 paschke Fan speed scan
  • 287005 08/30/09 05:05 pbaturin Fan speed set to 67Hz
  • 287003 08/30/09 04:57 pbaturin Fan Scan
  • 287002 08/30/09 04:55 paschke fan speed vs. cooling power
  • 287001 08/30/09 04:50 R. Michaels ADC18 Test Crate Study (Part II)
  • 286999 08/30/09 04:48 paschke Detector relay removed from L-HRS
  • 286996 08/30/09 04:28 R. Michaels ADC18 Test Crate study (Part I)
  • 286993 08/30/09 04:15 R. Michaels re: ADC18 in ROC26, 4th slot
  • 286989 08/30/09 03:55 A. Camsonne Compton electron detector run 18475
  • 286984 08/30/09 03:48 ahmed other target strips
  • 286983 08/30/09 03:47 z. Ahmed. L3 Power snap shot.
  • 286982 08/30/09 03:44 Z. Ahmed. L3 Temp snap shot.
  • 286981 08/30/09 03:43 Z. Ahmed. L3Power snap shot.
  • 286980 08/30/09 03:42 Z. Ahmed Main target screen.
  • 286978 08/30/09 03:36 pbaturin Use run 11751 instead of run 11739 for raster scan study
  • 286975 08/30/09 02:47 A. Camsonne First compton on Electron detector
  • 286971 08/30/09 02:24 M Friend Compton Photon Detector HV
  • 286970 08/30/09 02:18 adaq Magnet Cryogen Levels
  • 286964 08/30/09 02:09 Z. Ahmed BPM strip tool snap shot
  • 286962 08/30/09 02:07 Z. Ahmed Hall A Tools snap shot.
  • 286961 08/30/09 01:49 Z. Ahmed Controlled access
  • 286947 08/30/09 01:00 pbaturin Raster Scan Datasheet
  • 286945 08/30/09 00:52 paschke adc18 pedestal peculiarity
  • 286941 08/30/09 00:40 paschke 5x5 raster
  • 286936 08/30/09 00:31 M Friend Problem with Compton DAQ
  • 286933 08/30/09 00:05 dalton Lumi widths for MMC raster 7.7 x 8.0
  • 286932 08/30/09 00:03 R. Michaels more HRS scaler issues
  • 286929 08/30/09 00:01 Kiadtisak Swing Shift Summary
  • Saturday

  • 286924 08/29/09 23:43 dalton Lumi widths for MMC raster 7.7 x 5.6
  • 286920 08/29/09 23:39 dalton Lumi widths for MMC raster 4.4*4.0
  • 286919 08/29/09 23:37 Aniol target hph drops to zero
  • 286918 08/29/09 23:35 R. Michaels update scalers in online replay (HRS)
  • 286917 08/29/09 23:34 R. Michaels onlana.C repair
  • 286916 08/29/09 23:31 Kiadtisak Hall A Tools snapshot
  • 286913 08/29/09 23:19 Kiadtisak Spot++
  • 286910 08/29/09 23:10 R. Michaels L-HRS scalers in HRS datastream now
  • 286906 08/29/09 22:47 M Friend Compton Photon Detector HV Trip
  • 286904 08/29/09 22:01 paschke right arm detector relay fixed, removed helicity from RHRS
  • 286900 08/29/09 21:25 Kiadtisak Typos
  • 286899 08/29/09 21:14 Kiadtisak HAPPEX detector's footprint studies
  • 286898 08/29/09 21:10 Kiadtisak HAPPEX detector's footprint studies (Right arm)
  • 286895 08/29/09 20:38 Kiadtisak Image for my previous log
  • 286894 08/29/09 20:36 Kiadtisak HAPPEX detector's footprint studies (Left arm)
  • 286893 08/29/09 20:23 Kiadtisak HAPPEX detector's footprint studies (Left arm)
  • 286889 08/29/09 19:46 Aniol target screen
  • 286888 08/29/09 19:45 Aniol target temp graphs
  • 286871 08/29/09 18:15 paschke target update
  • 286866 08/29/09 18:03 aniol target - high power heater graph
  • 286861 08/29/09 17:39 mercado Wrong Figure3
  • 286860 08/29/09 17:38 mcnulty Larm Hapdet alignment
  • 286859 08/29/09 17:34 mercado Analysis of Lumi/Detector runs with Hydrogen
  • 286854 08/29/09 17:09 A. Camsonne Compton Electron detector current
  • 286853 08/29/09 17:06 A. Camsonne ETROC firmware version loaded from triggr_ctrl_restored
  • 286848 08/29/09 16:43 mcnulty Dummy Al spectra (regarding HapDet alignment)
  • 286844 08/29/09 16:12 mercado Target screenshot
  • 286843 08/29/09 16:11 mcnulty Detector Alignment tweak#1
  • 286842 08/29/09 16:09 mcnulty Proof of black box outline
  • 286841 08/29/09 16:08 buddhini Target Control Screen
  • 286840 08/29/09 16:06 mercado Day Shift Summary
  • 286838 08/29/09 15:26 mercado Controlled Access
  • 286837 08/29/09 15:13 A. Camsonne Compton analyzer helicity fixed
  • 286835 08/29/09 14:59 mcnulty VDC, s0, and hapdet HV changes
  • 286834 08/29/09 14:48 paschke target: is this the flow that matters?
  • 286833 08/29/09 14:44 M Friend Correct Compton Configuration
  • 286832 08/29/09 14:43 paschke temp excursion target
  • 286831 08/29/09 14:30 A. Camsonne Compton software recompiled
  • 286830 08/29/09 14:26 A. Camsonne Compton webcam network switch dead
  • 286829 08/29/09 13:53 buddhini heater
  • 286828 08/29/09 13:46 adaq Hall A tools
  • 286827 08/29/09 13:35 Jack Left Arm VDC threshold, jumper installed
  • 286826 08/29/09 13:19 R. Michaels turn off scripts for Test Crate (TestBob)
  • 286825 08/29/09 13:11 R. Michaels DAC noise off in CRL because its set by greenmonster
  • 286824 08/29/09 12:40 R. Michaels 18-bit developments in test cage
  • 286823 08/29/09 12:32 mcnulty Rarm alignment analysis update
  • 286818 08/29/09 11:57 R. Michaels triangle DAC noise default
  • 286813 08/29/09 11:15 R. Michaels Turned on DAC noise for ADC18
  • 286808 08/29/09 11:02 suhring ESR update
  • 286801 08/29/09 10:27 Rakitha (Hal typo in the trigger voltages
  • 286768 08/29/09 08:02 dalton New compton run
  • 286766 08/29/09 08:00 J. Wexler Shift Summary - Owl 8/29
  • 286763 08/29/09 07:19 dalton Feedback is on IA
  • 286760 08/29/09 06:58 dalton default.cuts changed
  • 286759 08/29/09 06:47 dalton Feedback turned on
  • 286758 08/29/09 06:43 dalton PITA slope determined for bcm 1
  • 286751 08/29/09 05:17 dalton BCM and happex detector pedestals calibrated
  • 286750 08/29/09 05:08 J. Wexler Changed L. Arm HAPPEX voltage
  • 286749 08/29/09 05:06 J. Wexler Shift Checklist - Owl Hall A Tools screen
  • 286746 08/29/09 05:02 J. Wexler Hall A Checklist - Owl Target Screen
  • 286743 08/29/09 04:43 paschke notes on 60 uA on LH2
  • 286742 08/29/09 04:42 J. Wexler Power and temperature of LH2 target shortly after beam trips
  • 286734 08/29/09 02:06 J. Wexler Switched to LH2 target
  • 286733 08/29/09 02:02 paschke spot++
  • 286732 08/29/09 01:54 J. Wexler Switch to BeO target.
  • 286695 08/29/09 01:02 Paul Baturin Target C hole
  • 286694 08/29/09 01:02 Paul Baturin PS8=10 changed to PS8=1
  • 286688 08/29/09 00:11 O.Glamazdin Moller results
  • Friday

  • 286686 08/28/09 23:55 Aniol shift summary
  • 286685 08/28/09 23:11 Kiadtisak HAPPEX detector's footprint studies (Right arm)
  • 286684 08/28/09 22:59 Kiadtisak HAPPEX detector's footprint studies (Right arm)
  • 286683 08/28/09 22:26 Kiadtisak HAPPEX detector's footprint studies (Left arm)
  • 286682 08/28/09 22:21 Kiadtisak HAPPEX detector's footprint studies (Left arm)
  • 286678 08/28/09 19:18 A. Camsonne ETROC photon external trigger firmware loaded
  • 286651 08/28/09 18:37 Aniol cryotarget stabilizes
  • 286650 08/28/09 18:34 Aniol stability of cryotarget
  • 286642 08/28/09 18:09 silwal Happex det linearity studies with led pulsers
  • 286638 08/28/09 17:57 R. Subedi HV ch for S2 and PVDIS
  • 286635 08/28/09 17:45 silwal DAQ check, after moving the startfiles into hallA system
  • 286623 08/28/09 17:13 mercado Lumi Runs on Carbon and Hydrogen
  • 286622 08/28/09 17:12 A. Camsonne Compton software recompiled in /home/compton/Aug-28-2009
  • 286620 08/28/09 17:09 paschke controlled access for FSD hardware, cavity electronics
  • 286616 08/28/09 16:40 R. Michaels DAQ back up and running. Qweak software LOCAL now.
  • 286601 08/28/09 16:02 Saha Shift Summary
  • 286600 08/28/09 16:02 aniol target screen
  • 286592 08/28/09 15:31 gaskell gzerol3 up
  • 286590 08/28/09 14:33 R. Michaels connectivity problems g0inj
  • 286586 08/28/09 14:18 A. Camsonne Access planned today at 4PM
  • 286579 08/28/09 13:04 paschke spect hv turned off (except for two special tubes)
  • 286578 08/28/09 13:04 Saha Hall A Tools
  • 286570 08/28/09 12:41 paschke BCM calibration runs, BPM calibration run
  • 286551 08/28/09 10:56 mercado Calibrating BCMs for 100uA
  • 286542 08/28/09 09:16 buddhini raster run 24587
  • 286537 08/28/09 08:56 buddhini raster run#24586
  • 286519 08/28/09 08:08 J. Wexler Shift summary - Owl 8/28
  • 286516 08/28/09 07:51 Reimer target empty
  • 286515 08/28/09 07:49 J. Wexler Harp scans and intrinsic size
  • 286514 08/28/09 07:41 Reimer target hydrogen
  • 286513 08/28/09 07:37 J. Wexler Raster map #2
  • 286508 08/28/09 07:19 J. Wexler Raster map #1
  • 286504 08/28/09 06:29 J. Wexler Carbon hole results and proper tune
  • 286497 08/28/09 06:09 Reimer raster++
  • 286492 08/28/09 06:00 Reimer Carbon hole
  • 286487 08/28/09 05:42 Reimer Target Carbon hole
  • 286486 08/28/09 05:38 J. Wexler Hall A Tools screen capture - Owl 8/28
  • 286483 08/28/09 05:33 paschke injector DAQ crate down
  • 286472 08/28/09 04:57 Reimer Target hydrogen
  • 286469 08/28/09 04:48 mcnulty Rarm Hapdet alignment study
  • 286463 08/28/09 04:14 A. Camsonne Compton Electron detector scope trace
  • 286457 08/28/09 04:04 mcnulty Larm HapDet alignment study
  • 286454 08/28/09 03:54 Reimer Target Thick Tantalum
  • 286451 08/28/09 03:48 J. Wexler Changed voltages on S0
  • 286448 08/28/09 03:42 A. Camsonne Compton timing
  • 286439 08/28/09 03:17 J. Wexler Sieve out.
  • 286438 08/28/09 03:16 silwal scaler rates on the counting daq
  • 286437 08/28/09 03:10 Reimer Target dummy
  • 286434 08/28/09 03:05 A. Camsonn Compton high current run
  • 286429 08/28/09 02:45 mcnulty Happex Detector Results from Hydrogen
  • 286428 08/28/09 02:31 A. Camsonne Owl shift run plan
  • 286425 08/28/09 02:18 Reimer Target Optics
  • 286414 08/28/09 01:55 J. Wexler Sieve in.
  • 286407 08/28/09 01:10 Reimer Target hydrogen
  • 286406 08/28/09 01:06 J. Wexler Position tune for target center
  • 286405 08/28/09 01:03 J. Wexler 2mm x 2mm raster
  • 286390 08/28/09 00:39 Reimer Target Motion OK
  • 286389 08/28/09 00:25 Reimer Target BeO
  • 286388 08/28/09 00:16 Reimer target hydrogen
  • 286385 08/28/09 00:13 J. Wexler Mirror missing from PREX detector in 8/25 tests
  • 286380 08/28/09 00:04 reimer Target screen
  • 286378 08/28/09 00:01 Xiaohui Shift Summary
  • Thursday

  • 286377 08/27/09 23:59 O.Glamazdin Moller result
  • 286376 08/27/09 23:57 Xinhu Yan target moved to optics
  • 286369 08/27/09 17:58 Xinhu Yan target IOC rebooted
  • 286368 08/27/09 17:57 Xinhu Yan target IOC rebooted
  • 286367 08/27/09 17:13 jpchen cryotarget is using 4K coolant, information for target operator
  • 286366 08/27/09 16:37 Xinhu Yan Target charts
  • 286365 08/27/09 16:27 Xiaohui Prepare for moller measurement
  • 286361 08/27/09 15:53 Higinbotham The Carbon Hole (x=-1.8mm,y=2.2mm)
  • 286336 08/27/09 15:21 Higinbotham Hole and No Hole
  • 286311 08/27/09 14:56 HAPPEX det linearity studies
  • 286310 08/27/09 14:56 HAPPEX det linearity studies
  • 286293 08/27/09 14:37 mcnulty Pulse height dist for Larm HAPdet
  • 286292 08/27/09 14:32 A. Camsonne Compton analog signal triggered by photon
  • 286281 08/27/09 14:15 Higinbotham Multifoil Run Complete
  • 286280 08/27/09 14:15 Reimer Target Carbon hole
  • 286276 08/27/09 13:44 Reimer Target Optics
  • 286272 08/27/09 13:16 mcnulty Attenuation coefficient for Rarm HAPdet
  • 286267 08/27/09 13:02 mcnulty Pulse height dist for Rarm HAPdet
  • 286262 08/27/09 11:50 Higinbotham Test During Compton Tuning
  • 286257 08/27/09 11:26 Higinbotham Found Left Arm Sieve In
  • 286254 08/27/09 11:05 Higinbotham pi [MeV/c]
  • 286253 08/27/09 11:04 Jack Left Arm VDC threshold
  • 286250 08/27/09 10:50 Reimer bta
  • 286247 08/27/09 10:14 Techs/spiege Reset Magnets
  • 286246 08/27/09 09:58 meekins Cryotarget support from ESR.
  • 286245 08/27/09 09:57 Higinbotham Cryogenetic Problems and Access
  • 286244 08/27/09 09:40 V. Sulkosky Re: HRS VDC status
  • 286242 08/27/09 09:33 reimer target cryoflow
  • 286238 08/27/09 08:58 A. Camsonne Compton GSO photon detector in beam position for background reduction tune
  • 286225 08/27/09 08:00 J. Wexler Shift Summary - Owl 8/27
  • 286222 08/27/09 07:27 A. Camsonne Electron detector noise snapshot
  • 286217 08/27/09 07:12 A. Camsonne Compton GSO crystal HV on
  • 286207 08/27/09 06:15 silwal Switching between Integrating and Counting DAQ (on happex det)
  • 286200 08/27/09 05:59 A. Camsonne One new screen put on the controls screen computer
  • 286196 08/27/09 05:53 A. Camsonne BeO check
  • 286185 08/27/09 05:28 J. Wexler Beam and magnets
  • 286178 08/27/09 04:46 A. Camsonne Backing out from Moller tune - getting ready for Lumi and HAPPEX detector tests
  • 286177 08/27/09 04:02 A. Camsonne Changed scripts to output to /adaql10/work1/happexsp
  • 286176 08/27/09 03:57 A. Camsonne Left arm cosmics check-out.
  • 286175 08/27/09 03:53 A. Camsonne HRS VDC status
  • 286174 08/27/09 03:35 O.Glamazdin Moller problems
  • 286171 08/27/09 03:26 A. Camsonne Disk clean-up
  • 286170 08/27/09 03:04 J. Wexler Controlled access
  • Wednesday

  • 286166 08/26/09 22:19 K. Kumar controlled access
  • 286161 08/26/09 19:01 Jack (Techs) Magnet power supplies
  • 286160 08/26/09 18:56 M Friend Before Compton Tune: Compton Photon Detector
  • 286159 08/26/09 18:39 Reimer access to reset quads
  • 286158 08/26/09 18:27 Reimer Max Current Test 70 muA
  • 286154 08/26/09 18:16 Reimer attempting 100muA
  • 286152 08/26/09 17:50 K. Kumar Immediate plan
  • 286151 08/26/09 17:28 Reimer High Current Running
  • 286150 08/26/09 17:06 reimer ESR and Accelerator Status
  • 286139 08/26/09 16:09 Higinbotham Target Gone Too
  • 286136 08/26/09 16:04 Higinbotham Shift Summary
  • 286135 08/26/09 16:03 A. Camsonne Left HRS HV back on and trigger downloaded
  • 286134 08/26/09 15:59 atrig Trigger-Download
  • 286133 08/26/09 15:43 J.H.Lee ESR crash
  • 286132 08/26/09 15:34 Higinbotham ESR Trip
  • 286130 08/26/09 15:31 atrig Trigger-Download
  • 286121 08/26/09 14:32 Higinbotham Momentum Set To pi [MeV/c]
  • 286120 08/26/09 12:32 R. Subedi access activity
  • 286108 08/26/09 09:37 dalton dalton
  • 286107 08/26/09 09:36 dalton RF recovery on Thursday is CANCELLED
  • 286102 08/26/09 09:18 mcnulty 45deg slanted quartz det results
  • 286098 08/26/09 08:34 Jeong Han Le Target thick tantalum
  • 286095 08/26/09 08:00 J. Wexler Shift summary 8/26 owl
  • 286091 08/26/09 07:47 J. Wexler Run checklist screen capture - 8/26 Owl
  • 286087 08/26/09 06:59 Reimer Target thin tantalum
  • 286086 08/26/09 06:58 J. Wexler PREX detector change
  • 286085 08/26/09 06:46 R. Michaels switch to HapNoL
  • 286080 08/26/09 06:19 R. Michaels Lumi runs with "high" current
  • 286064 08/26/09 05:39 J. Wexler MCC unable to reach requested current
  • 286062 08/26/09 05:30 R. Michaels ion chamber update
  • 286060 08/26/09 05:08 R. Michaels update -- ion chamber work
  • 286059 08/26/09 04:26 R. Michaels more runs. higher current, ovs sturdy
  • 286056 08/26/09 04:20 Reimer Target Thick Tantalum
  • 286053 08/26/09 04:07 mcnulty preliminary Prex det pulse-height dists
  • 286051 08/26/09 03:58 R. Michaels happex_test.crl change
  • 286049 08/26/09 03:56 R. Michaels DAQ study of oversampling
  • 286044 08/26/09 03:43 R. Michaels some oversample runs
  • 286035 08/26/09 03:28 Reimer Target on Empty
  • 286033 08/26/09 03:20 Reimer Target thick tantalum
  • 286031 08/26/09 03:17 Re: run 11631 data changes
  • 286029 08/26/09 03:06 R. Michaels restore happex_test.crl
  • 286028 08/26/09 03:04 R. Michaels HAPPEX DAQ: Some problems resolved.
  • 286027 08/26/09 03:04 J. Wexler MCC acknowledges current plan requests
  • 286026 08/26/09 02:57 Reimer Target empty
  • 286024 08/26/09 02:53 R. Michaels strange things in HAPPEX analysis
  • 286023 08/26/09 02:42 J. Wexler Current ramp program requested
  • 286020 08/26/09 02:38 R. Michaels HAPPEX DAQ startup, part I
  • 286006 08/26/09 02:00 Reimer Target thick tantalum
  • 285995 08/26/09 01:10 A. Camsonne HV off on detectors and VDC on left arm
  • 285991 08/26/09 01:01 J. Wexler PREX magnet tune
  • 285986 08/26/09 00:32 Reimer Target Thin Tantalum
  • 285985 08/26/09 00:15 J. Wexler Current PREX configuration
  • 285984 08/26/09 00:09 Reimer target screen
  • 285983 08/26/09 00:02 Xiaohui shift summary
  • Tuesday

  • 285981 08/25/09 23:37 A. Camsonne Tune beam in Hall
  • 285980 08/25/09 23:37 Navaphon JT Valve reduce loop 3
  • 285979 08/25/09 23:23 A. Camsonne GSO HV tripped
  • 285978 08/25/09 23:10 A. Camsonne (Techs) Bottom VDC Left arm discriminator read back low
  • 285973 08/25/09 21:27 D. Parno Compton FADC: Helicity-Correlated Pedestal Shift
  • 285972 08/25/09 19:46 Xiaohui update
  • 285971 08/25/09 18:49 dalton Hall status
  • 285970 08/25/09 18:47 Navaphon Target position
  • 285969 08/25/09 18:46 A. Camsonne Compton electron detector noise study and tests
  • 285968 08/25/09 18:43 A. Camsonne Compton electron detector noise study and tests
  • 285967 08/25/09 18:31 R. Michaels Quartz HV on
  • 285965 08/25/09 18:03 R. Michaels Oversampling fixed (I think).
  • 285958 08/25/09 17:57 jpchen cryotarget cooled-down with Loop 3 LH2.
  • 285957 08/25/09 17:55 J. Wexler PREX detector in place
  • 285956 08/25/09 17:52 mercado Sync Signals from CH and R-RHS for OS=1,2,4, runs 11556, 11557, 11558
  • 285943 08/25/09 17:32 D. Parno Compton: FADC pedestal value for GSO Channel
  • 285935 08/25/09 17:23 R. Michaels Quartz pulse-height spectrum, cosmics
  • 285924 08/25/09 16:41 Bob/KK prex quartz det work, part II
  • 285918 08/25/09 16:33 Xiaohui MCC update
  • 285912 08/25/09 16:14 Bob/KK quartz det work, part I
  • 285910 08/25/09 16:12 Higinbotham Shift Summary
  • 285899 08/25/09 15:37 rsholmes Timeslots are okay
  • 285896 08/25/09 14:54 D. Parno Compton: FADC Accumulator Sum Rule
  • 285891 08/25/09 14:04 rsholmes pan errors
  • 285890 08/25/09 14:03 rsholmes pan errors
  • 285886 08/25/09 13:26 Higinbotham Magnet Ramp Down
  • 285884 08/25/09 12:43 R. Michaels Oversampling - working, sorta
  • 285883 08/25/09 12:31 dalton Accelerator energy change until 6 pm
  • 285881 08/25/09 12:22 Jack (Techs) Left Shield House air conditioner
  • 285879 08/25/09 12:14 Jack (Techs) Left Dipole serial cable
  • 285874 08/25/09 11:52 M Friend Dithering Signal in Compton DAQ
  • 285859 08/25/09 11:30 Moffit BUG found in flexio Library...
  • 285838 08/25/09 09:52 dalton Beam energy measurement
  • 285837 08/25/09 09:28 mcnulty Prex test runlist
  • 285834 08/25/09 08:00 J. Wexler Shift summary - owl 8/25/2009
  • 285833 08/25/09 07:53 mercado run numbers for Thick Tantalum 5.5uA on HAPPEX DAQ
  • 285832 08/25/09 07:49 mercado Regressed asymmetries for Thick Tantalum target at ~5.5uA
  • 285831 08/25/09 07:34 mercado BCM3 moved to ADC23_3
  • 285817 08/25/09 06:56 J. Wexler 8-25 owl shift checklist - Hall A Tools capture
  • 285809 08/25/09 06:32 J. Wexler Switched T3 trigger in the right arm
  • 285808 08/25/09 06:31 Reimer Target to thick tantalum
  • 285797 08/25/09 04:45 J. Wexler 45 deg. PREX detector mechanical malfunction
  • 285796 08/25/09 04:39 R. Michaels Notes on access
  • 285795 08/25/09 04:36 R. Michaels oversampling not working
  • 285784 08/25/09 03:34 A. Camsonne Left arm crates off
  • 285776 08/25/09 02:53 Reimer controlled access
  • 285757 08/25/09 02:15 R. Michaels HapNoL mod
  • 285747 08/25/09 01:39 J. Wexler PREX magnet tune set
  • 285746 08/25/09 01:19 Reimer Target at Thin Tantalum
  • 285745 08/25/09 00:58 J. Wexler Compton ion chamber issue
  • 285744 08/25/09 00:04 Xiaohui Shift summary
  • Monday

  • 285738 08/24/09 22:13 A. Camsonne HV off on left arm - Left hut at 88.5 degrees
  • 285726 08/24/09 20:44 M Friend Change to Compton Clock
  • 285725 08/24/09 20:23 A. Camsonne Module temperature reasonable for now
  • 285724 08/24/09 20:08 R. Michaels new HapNoL config
  • 285722 08/24/09 19:20 dalton Left arm cooling update
  • 285721 08/24/09 19:16 Xiaohui Beam permit
  • 285719 08/24/09 19:08 Navaphon target moves to carbon hole
  • 285718 08/24/09 18:31 A. Camsonne Etroc output cables to upstair patch
  • 285716 08/24/09 18:18 Jack Right Arm VDC HV power supply
  • 285715 08/24/09 18:15 Jack (Techs) Left Arm Shield House cooling
  • 285712 08/24/09 16:35 Higinbotham Shift Summary
  • 285708 08/24/09 15:49 mercado Regressed Asymmetry results for Run 11547 - Tantalum at 20uA
  • 285707 08/24/09 15:38 Jack Right Arm VDCs
  • 285699 08/24/09 13:08 M Friend Compton Wiring Changes
  • 285693 08/24/09 11:48 mercado More Lumi commisioning. Attempt at Oversampling.
  • 285692 08/24/09 11:40 dalton Epics reboot reestablishes compton laser slow ramp
  • 285688 08/24/09 11:17 fansler RQ2 Return
  • 285683 08/24/09 10:52 dalton New metrolab available for 9th dipole map
  • 285680 08/24/09 09:58 R. Subedi T3 fixed
  • 285678 08/24/09 08:57 Higinbotham Setting Spectrometers to Zero
  • 285668 08/24/09 08:00 mcnulty Changed channel for PREx Quartz detector
  • 285666 08/24/09 07:55 mcnulty RHRS T3 trigger missing?
  • 285665 08/24/09 07:52 mcnulty RHRS coda rates strange
  • 285663 08/24/09 07:46 mcnulty Start of PREx tune
  • 285657 08/24/09 07:00 Kai important Run #4298 for prex test
  • 285649 08/24/09 06:13 adaq S0 voltages
  • 285648 08/24/09 06:11 paschke additional optics study with sieve, watercell
  • 285632 08/24/09 04:38 paschke activity during access
  • 285631 08/24/09 04:36 paschke improved carbon spectrum in left arm after S0 swap
  • 285621 08/24/09 01:50 R. Michaels Run 24498 X-Y anomoly
  • 285620 08/24/09 01:27 paschke Right/Left HRS triggers at -2%
  • 285619 08/24/09 01:15 paschke dp plots from optics studies (corrected)
  • 285618 08/24/09 01:01 paschke focal plane distribution in left arm
  • 285613 08/24/09 00:07 paschke dp plots for right vs. left HRS
  • 285612 08/24/09 00:05 D. Parno Swing Shift Summary 8/23
  • Sunday

  • 285609 08/23/09 23:47 paschke focal plane y for Ta elastic
  • 285607 08/23/09 23:40 paschke contour plot of left arm focal plane
  • 285606 08/23/09 23:36 paschke right arm focal plane distributions
  • 285605 08/23/09 23:26 D. Parno Called for controlled access
  • 285604 08/23/09 23:25 paschke left arm trigger?
  • 285599 08/23/09 22:59 D. Parno HARP scan at (+2.5, +2.5)
  • 285576 08/23/09 22:00 D. Parno Moved to carbon target for bull's-eye scan
  • 285570 08/23/09 21:41 D. Parno Beam is back
  • 285567 08/23/09 21:18 Kiadtisak Full analysis for optics studies
  • 285566 08/23/09 21:13 D. Parno Moved to water target
  • 285559 08/23/09 20:57 D. Parno Moved to carbon target for beam return
  • 285555 08/23/09 20:37 D. Parno Right VDC Level Dropping
  • 285552 08/23/09 20:10 J. Wexler Sieve is in place.
  • 285551 08/23/09 20:10 D. Parno Beam down to fix shunt
  • 285546 08/23/09 19:47 D. Parno Moved to water target
  • 285541 08/23/09 19:35 Apar Regressed asymmetries for Tantalum at ~10uA
  • 285528 08/23/09 19:25 J. Wexler Run information incorrect on 24490-93, 4275-8
  • 285527 08/23/09 19:22 riordan Right VDC Discriminator History
  • 285526 08/23/09 19:04 D. Parno Moved to carbon target
  • 285517 08/23/09 18:00 D. Parno Sieve out, tantalum target in
  • 285516 08/23/09 17:51 R. Michaels opinion re: prescale factors
  • 285515 08/23/09 17:40 D. Parno Ion chamber problems; moved to empty target for beam tune
  • 285510 08/23/09 17:20 D. Parno Leaving water target at 5 uA for now
  • 285509 08/23/09 17:17 A. Camsonne Added VDC efficiencies in the online displays
  • 285508 08/23/09 17:13 D. Parno Moved to water target
  • 285507 08/23/09 17:12 A. Camsonne rootlogon script modified to load online display
  • 285506 08/23/09 17:07 A. Camsonne Dead wires in Right Arm VDC
  • 285503 08/23/09 17:02 J. Wexler VDC Disc. Level
  • 285497 08/23/09 16:25 A. Camsonne moved HV server to adaql5
  • 285493 08/23/09 16:09 Kai Pan Day shift summay
  • 285492 08/23/09 16:08 A. Camsonne High Voltage GUI for right arm
  • 285490 08/23/09 15:57 M Friend Compton High Voltage Channels
  • 285479 08/23/09 15:35 Kai target move to Ta
  • 285476 08/23/09 15:31 Kai spot++ run #24488
  • 285473 08/23/09 15:29 Kiadtisak The image of the plot for run# 24481
  • 285472 08/23/09 15:09 Kiadtisak Run#24481 plot (momentum with cut)
  • 285471 08/23/09 15:07 R Michaels
  • 285470 08/23/09 15:04 paschke partiall run list for optics studies
  • 285469 08/23/09 15:01 Kiadtisak Run#24481 (tg_y and th:ph)
  • 285468 08/23/09 14:28 paschke more on trigger cuts in analyzing spectrometer data
  • 285467 08/23/09 14:28 adaq spot++ run#24487
  • 285466 08/23/09 14:08 R. Michaels PREX Quartz off or unplugged
  • 285465 08/23/09 14:07 Kiadtisak Run #4244 plot (th vs ph and momentum)
  • 285464 08/23/09 13:52 Kai prescale changed
  • 285459 08/23/09 13:42 Apar pan problem fixed
  • 285455 08/23/09 13:32 dalton Compton locking without slow ramp
  • 285449 08/23/09 13:13 Kiadtisak Right arm's plots (tg_y and tg_th:tg_ph) for watercell
  • 285447 08/23/09 13:06 R. Michaels spectra for run 4258
  • 285445 08/23/09 13:02 Kai Pan change p0 to 1.1660, -3.0%
  • 285440 08/23/09 12:50 R. Michaels re: D.evtypebits
  • 285439 08/23/09 12:38 Kiadtisak Left arm's plot (momentum for watercell target) corrected
  • 285438 08/23/09 12:37 paschke trigger fine
  • 285437 08/23/09 12:35 R. Michaels correlation to HAPPEX det
  • 285436 08/23/09 12:34 Kiadtisak Left arm (tg_y and th:ph) Corrected
  • 285435 08/23/09 12:34 paschke Left arm trigger broken
  • 285422 08/23/09 11:46 mcnulty HV mapping for HAPPEX
  • 285419 08/23/09 11:36 R. Michaels dP/P for water
  • 285418 08/23/09 11:32 Kiadtisak Right arm plot for tg_y and tg_th:tg_ph
  • 285415 08/23/09 11:24 R. Michaels R-HRS sieve run 4247
  • 285414 08/23/09 11:23 Kiadtisak momentum plot after cutting for watercell
  • 285411 08/23/09 11:15 R. Michaels data checks
  • 285410 08/23/09 11:13 Kiadtisak Plot tg_y and tg_th:tg:ph
  • 285407 08/23/09 11:01 Kai Pan making up halog about target movement
  • 285402 08/23/09 10:31 Kai Pan change to "sieve in"
  • 285397 08/23/09 10:03 Kai current sieve status: OUT
  • 285396 08/23/09 09:49 Kai Pan spot++ L#24474 (R # 4256 doesn't work)
  • 285383 08/23/09 08:53 kiadtisak carbon optics sieve out pattern
  • 285371 08/23/09 08:02 xqian Owl Shift Summary
  • 285327 08/23/09 05:02 adaq Run 24455 info corrected
  • 285323 08/23/09 04:46 Kai Pan carbon optics sieve in pattern
  • 285314 08/23/09 04:00 paschke pan analysis issues
  • 285303 08/23/09 02:59 adaq sieve pattern for H20
  • 285296 08/23/09 02:30 paschke plot of VDC current
  • 285295 08/23/09 02:28 paschke VDC currents still unstable
  • 285294 08/23/09 02:26 adaq dp plot for H20
  • 285289 08/23/09 02:08 paschke dp for carbon, +1.5%, no raster
  • 285278 08/23/09 01:42 xqian dp spectrum on Tantalum
  • 285274 08/23/09 01:34 paschke position, energy strips for early optics running
  • 285273 08/23/09 01:32 paschke RHRS VDC current reading unstable?
  • 285270 08/23/09 01:30 adaq delta_p cutting the central foils on right arm
  • 285269 08/23/09 01:29 mercado Lumi HVs for Tantalum at 8.5uA
  • 285268 08/23/09 01:27 xqian delta_p cutting on the central foils
  • 285265 08/23/09 01:15 mercado Tuned HVs for Carbon optics at 5uA.
  • 285257 08/23/09 00:57 mcnulty spot++ for 2.5mm raster
  • 285251 08/23/09 00:47 mcnulty spot++ for optics calibration
  • 285248 08/23/09 00:44 K. Kumar Raster calibration
  • Saturday

  • 285225 08/22/09 23:40 K. Kumar, J. Swing shift summary
  • 285224 08/22/09 23:36 R. Michaels Order restored to scalers. Map of channels
  • 285221 08/22/09 23:21 R. Michaels DAQ restored, ROC2 replaced.
  • 285220 08/22/09 23:08 K. Kumar Access work completed
  • 285216 08/22/09 21:43 K. Kumar Request controlled access
  • 285214 08/22/09 21:39 K. Kumar Position of beam
  • 285212 08/22/09 21:32 R. Michaels R-HRS DAQ unstable. Access needed.
  • 285211 08/22/09 21:27 R. Michaels things not making sense in scalers
  • 285209 08/22/09 21:20 R. Michaels HRS DAQ ROC2 problems.
  • 285207 08/22/09 21:07 K. Kumar 7/18 harp scans
  • 285204 08/22/09 21:02 K. Kumar 3a/3b harp scan
  • 285203 08/22/09 20:01 K. Kumar Beam finally
  • 285202 08/22/09 19:44 K. Kumar, J. Collimators moved
  • 285201 08/22/09 19:10 K. Kumar Collimator settings
  • 285200 08/22/09 18:57 K. Kumar Beam in hall finally
  • 285196 08/22/09 16:04 Kai Pan Day Shift Summary
  • 285195 08/22/09 15:23 gomez NOTE: NEW HRSR & HRSL COLLIMATOR POSITIONS
  • 285194 08/22/09 13:07 dalton Compton slow ramp not working
  • 285192 08/22/09 12:03 R. Michaels timebrd.cfg updating-mechanism solution
  • 285185 08/22/09 09:48 dalton Target list
  • 285184 08/22/09 09:41 dalton Beam update
  • 285182 08/22/09 07:56 xqian Owl Shift Summary
  • Friday

  • 285180 08/21/09 23:16 K. Kumar Swing shift summary
  • 285179 08/21/09 22:45 A. Camsonne Locking cavity after EPICS reboot painful
  • 285178 08/21/09 22:10 K. Kumar Another beam update
  • 285177 08/21/09 22:06 A. Camsonne Rebooted Comption EPICS IOC
  • 285176 08/21/09 21:43 R. Subedi Happex signals in Scalers
  • 285175 08/21/09 21:32 Ramesh/Dusti HRS triggers ready
  • 285174 08/21/09 21:20 paschke detector hut configuration
  • 285173 08/21/09 20:57 R. Michaels re: ROC2
  • 285172 08/21/09 20:06 K. Kumar Beam update
  • 285170 08/21/09 19:32 K. Kumar Beam still down
  • 285168 08/21/09 19:21 A. Camsonne Warning message for ROC2
  • 285167 08/21/09 19:21 A. Camsonne Warning message for ROC2
  • 285165 08/21/09 18:49 A. Camsonne Compton parameter change for software matching electron and photon
  • 285164 08/21/09 18:46 dalton (for PREX-like tune for HRS
  • 285160 08/21/09 18:34 dalton Day summmary
  • 285159 08/21/09 18:01 dalton Hall in controlled access
  • 285158 08/21/09 16:31 R. Subedi RHRS S0 TDC ok
  • 285154 08/21/09 14:53 mercado Pedestal noise for different 18-bit ADC gain settings
  • 285152 08/21/09 14:36 atrig Trigger-Download
  • 285151 08/21/09 14:27 atrig Trigger-Download
  • 285150 08/21/09 14:26 atrig Trigger-Download
  • 285149 08/21/09 14:20 R. Michaels RT unplugged (not needed)
  • 285148 08/21/09 14:16 atrig Trigger-Download
  • 285146 08/21/09 14:15 atrig Trigger-Download
  • 285145 08/21/09 14:14 atrig Trigger-Download
  • 285144 08/21/09 14:09 atrig Trigger-Download
  • 285143 08/21/09 14:08 atrig Trigger-Download
  • 285141 08/21/09 13:21 folts tech
  • 285140 08/21/09 13:18 pkpan FADC turned off on both arms
  • 285135 08/21/09 12:43 D. Parno VtoF Card #3 in Compton DAQ
  • 285132 08/21/09 12:32 Jack (Techs) Right Spectrometer Dipole work
  • 285130 08/21/09 12:14 R. Subedi hall work this morning
  • 285129 08/21/09 12:03 D. Parno VtoF Card #2 in Compton DAQ
  • 285100 08/21/09 10:04 mercado Pedestals
  • 285096 08/21/09 09:52 dalton Lock hall at noon
  • 285095 08/21/09 09:46 mercado Pedestal values for Lumis and HAPPEX detectors
  • 285092 08/21/09 07:56 xqian Owl Shift Summary
  • 285091 08/21/09 07:47 paschke Hall A tools (with beam)
  • 285090 08/21/09 07:39 paschke Hall A tools
  • Thursday

  • 285078 08/20/09 22:37 dalton BCMs added to CH parity DAQ
  • 285077 08/20/09 22:37 dalton BCMs added to CH parity DAQ
  • 285076 08/20/09 22:32 silwal correction on the y-axis: y-axis is the baseline ADC channels, not diff on - diff off
  • 285075 08/20/09 22:22 silwal LED tests: baseline intensity variation with frequency
  • 285073 08/20/09 21:36 D. Parno FADC Compton wiring: VtoF signals
  • 285070 08/20/09 20:17 dalton CW beam tonight
  • 285062 08/20/09 17:36 R Michaels HAPPEx DAQ status
  • 285047 08/20/09 16:15 Bob/Kai FADC on R-HRS, but disabled.
  • 285040 08/20/09 14:56 Techs / Spie Pre-Run Checklist
  • 285028 08/20/09 12:31 dalton Pulsed beam tonight
  • 285013 08/20/09 09:12 R. Michaels FADC in sfi2.crl
  • 285012 08/20/09 08:39 Bob/Kai/Rame PVDIS crate timing issue resolved
  • 285002 08/20/09 00:07 silwal relay switch test on the left arm parity daq, pedestal probably wrong!!
  • Wednesday

  • 284980 08/19/09 13:34 Bob/Zafar cavities in HRS DAQ
  • 284979 08/19/09 13:28 R. Michaels scaler cabling fixed
  • 284953 08/19/09 11:07 A. Camsonne Compton Electron detector FSD checked out
  • 284949 08/19/09 10:44 R. Michaels HAPPEX DAQ fallen, now revived
  • 284942 08/19/09 10:08 gunning raster
  • 284941 08/19/09 09:36 dalton New location for ROOTfiles
  • 284940 08/19/09 08:57 R. Subedi Happex Triggers, signals, and HV mapping in HRSs
  • 284939 08/19/09 08:56 R. Subedi Happex Triggers, signals, and HV mapping in HRSs
  • 284937 08/19/09 07:38 R. Michaels why boot from adev ?
  • 284935 08/19/09 00:18 paschke analysis errors fixed, pedestal run started
  • 284932 08/19/09 00:11 paschke testing left arm switch
  • Tuesday

  • 284930 08/18/09 23:17 paschke errors on ADCX unpack
  • 284926 08/18/09 23:13 paschke cable pulled from R-HRS parity daq
  • 284925 08/18/09 23:10 paschke, sil mapping the lumis tonight
  • 284924 08/18/09 21:19 paschke shouldn't the ROCs run as apar?
  • 284779 08/18/09 10:24 R. Michaels SYNC errors in rcServer resolved
  • Monday

  • 284626 08/17/09 09:45 R. Michaels HAPPEX DAQ for Happex and HapNoI
  • 284625 08/17/09 09:40 R. Michaels HAPPEX DAQ restored
  • Sunday

  • 284617 08/16/09 21:46 A. Camsonne ETROC switched to external clock
  • 284612 08/16/09 18:26 R. Michaels HAPPEx DAQ status : busted
  • 284590 08/16/09 12:37 R. Michaels devices update
  • Saturday

  • 284564 08/15/09 18:27 Jack (Techs) Left Shield House right side door
  • 284563 08/15/09 18:15 Jack VDC connections
  • 284562 08/15/09 18:15 R. Michaels re: DT with FADC
  • 284555 08/15/09 14:32 R.Michaels/K ROC4 swap, FADC readout update
  • Friday

  • 284427 08/14/09 12:50 dalton Started compton
  • 284351 08/14/09 10:01 A. Camsonne Compton Electron detector chamber closed and shielding in place
  • Thursday

    
    

    Wednesday

    
    

    Tuesday

    
    

    Monday

  • 283451 08/10/09 09:32 A. Camsonne Compton laser power supply locked and tagged
  • 283449 08/10/09 09:02 A. Camsonne Laser of on Compton cavity for beamline work
  • Sunday

    
    

    Saturday

    
    

    Friday

    
    

    Thursday

  • 0 08/06/09 15:58 no_name no_key
  • 0 08/06/09 15:58 no_name no_key
  • 0 08/06/09 15:58 no_name no_key
  • Wednesday

    
    

    Tuesday

  • 283204 08/04/09 18:47 A. Camsonne hatsv4 portserver acting odd
  • 0 08/04/09 16:03 no_name no_key
  • Monday

    
    

    Sunday

    
    

    Saturday

  • 283001 08/01/09 15:12 techs magnets
  • 283000 08/01/09 12:38 folts Both Q2s on cold return
  • 282999 08/01/09 12:28 folts Both Dipoles on cold return